OpenStudy (anonymous):

The speed that an airplane travels is calculated by the formula: g = a - h Ground speed (g) = air speed (a) minus headwind speed (h). Find the ground speed if the air speed = 305 mph and the headwind speed = 41 mph. A) 264 mph B) 346 mph C) 246 mph D) 364 mph

jimthompson5910 (jim_thompson5910):

g = a - h g = 305 - 41 g = ???

OpenStudy (anonymous):

264?

OpenStudy (anonymous):

so it would be A?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks!!!

jimthompson5910 (jim_thompson5910):

yw